Circular No. 6501

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1655–40 B. A. Harmon, C. R. Robinson, G. J. Fishman, S. N. Zhang, and W. S.

Paciesas, report for the Compton Observatory BATSE Team: “Hard x-ray flux from the superluminal jet source GRO J1655–40 has been increasing since mid-October, and has approximately doubled in intensity since Oct. 26, with fluxes near 1.0 Crab on Oct. 30–31 (20–100 keV). This activity follows an extended, bright, and highly-variable outburst (IAUC 6436) between June and Sept. Spectra obtained between Oct. 26 and 30 are adequately fit with a power law between 20 and 200 keV, with photon index changing from −2.6 to −2.9 (±0.1) as the source brightened.”

SGR 1806–20 C. Kouveliotou, Universities Space Research Association; G. J. Fish-

man, C. A. Meegan, Marshall Space Flight Center, NASA; J. van Paradijs, University of Amsterdam and University of Alabama in Huntsville (UAH); M. S. Briggs and G. Richardson, UAH; and K. Hurley, University of California, Berkeley, report: “The BATSE has triggered on renewed activity from the soft γ-ray repeater SGR 1806–20 on Oct. 30.95220 and on Oct. 31.54094 and 31.97024 UT. The fluence and peak flux of the first event are 10^{−7} erg cm^{−2} and 10^{−6} erg cm^{−2} s^{−1}, respectively. The reactivation of the source has initiated several Target of Opportunity observations with other satellites and groundbased observatories; currently RXTE has detected several bursts from the source and is observing during Nov. 6. We have changed the BATSE trigger criteria to trigger on (weaker) bursts during the RXTE observations. We strongly encourage follow-up observations in other wavelengths.”

COMET C/1996 Q1 (TABUR) Total magnitude estimates (visual unless otherwise noted): Oct. 31.76

UT, 8.7 (H. Mikuž, Ljubljana, Slovenia, 0.20-m f/2 Baker-Schmidt camera + V filter + CCD); Nov. 1.80, 8.3 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); 1.98, 9.5 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 2.75, 8.7 (Mikuž); 2.75, 9.2 (B. H. Granslo, near Tryvann, Oslo, Norway, 0.20-m reflector); 3.99, 9.5 (Bortle); 4.74, 8.6 (V. Znojil, Brno, Czech Republic, 25x100 binoculars).

1996 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green